Technical Observing: 30 December 2012 UT

Observer: JHill (LBTO)
Telescope Operator: GBechetti (LBTO)
Software Support: TSargent (Tucson)
AO Support: JChristou (Tucson)
Telescope Support: KNewton (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

We were open for ~2.5 hours doing various binocular tests. Eventually the clouds and wind forced us to close.

Details

Open

Clear sky in afternoon, but clouds and wind are increasing by sunset.

02:05 Clouds are thinning a bit, and winds are still slightly below 20 m/s, so we are opening

Authorize LUCI @ LFBG monocular

02:15 ACQUIRE Preset IE=-36 CA=-9

02:18 ACTIVE Preset BS9188 - seeing is 1.5 - 2 arcsec.

Binocular Tests

Authorize IRTC @ RFBG along with LUCI @ LFBG

Apply the RZ +5.12 deg global offset to DX M3

02:29 Left Async ACTIVE preset

02:31 Right ACQUIRE, then Left ACQUIRE

02:38 spiral both sides to move mount

then paddle to center each side. SX IE=-37.8 CA=-8.4 DX IE=47.46 CA=-66.76

Offset 0 -30 on left - this makes copointing warnings from both sides (one warning would be sufficient).

Issue: DX side is having problems with collimation. It looks like M2 is not putting any astigmatism LUT, and the DX primary has saturated at -20000 Z6 and is also applying -17000 Z5.

We called Julian to see if he can turn on the LUT. DX shell RIPs.

SX complaining about guide star not stable after 2 iterations. See Issue 4434.

02:47 Async ACTIVE presets each side

03:09 offset 0 0

03:10 Right shell RIPs

Julian restarts the arbitrator and everything.

03:45 Corrected the guide correction factor and guidecam_clipping and guideWindowSize in LUCI_IRTC_R.cfg

03:55 Resend right side preset after restarting GCSR. Right guide box parameters are now correct, but guide box as drawn on the acq image is still different (it took another preset to be redrawn?).

04:xx DX M2 hits +RY limit. Adjust Xglobal offset to be more negative. Xglobal=-1.5 but that is still not enough.

04:15 Xglobal to -1.75

04:20 SYNC ACTIVE presets to BS9107 - but no star? Geno's getting good practice with binocular spiralling and acquisition.

LUCI_L.cfg: guidecam_ROI "16 40 536 508"
LUCI_IRTC_R.cfg: guidecam_ROI "18 16 526 510"
LUCI_IRTC_L.cfg: guidecam_ROI "16 40 536 508"

Close

04:47 Closing as the clouds are getting seriously thick.

Closed Dome Binocular Testing

05:34 SYNC ACTIVE presets

05:38 Turn on fake guide stars in gcs.conf. Restart GCSs. Working on Issue 4431.

05:40 SYNC ACTIVE presets

05:48 Took out the guidecam_ROI parameters to see if they were blocking the drifting hotspot. Restart GCSs.

Aha! Not guidecam_ROI, but guidecam_clipping is the problem.

05:56 Took out guidecam_clipping. Restart GCSs.

05:58 SYNC GUIDE presets

Yes, the drifting hotspot blockage is caused by the guidecam_clipping parameter. Added note to Issue 4431.

-- JohnHill - 30 Dec 2012
Topic revision: r4 - 30 Dec 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback